# Articulatory Encodec: Coding Speech through Vocal Tract Kinematics

### v2

@@ VI Conclusion
We propose a novel encoding-decoding framework of speech, Articulatory Encodec, which mimics the biophysical interface of speech production. Through large-scale training, our proposed Articulatory Encodec framework achieves a high-performance articulatory inversion and synthesis that can generalize to unseen speakers, with a minimal loss of information compared to ground truth speech. To our knowledge, this is the first demonstration of universal articulatory synthesis that can scale up to an indefinite number of speakers. This universality is supported by a novel speaker identity encoding that embeds highly discriminable speaker-specific voice textures. Additionally, the encoded speaker embedding is effectively disentangled from articulatory features and can be used for accent-preserving voice conversion. Using the articulatory encodec, speech can be represented in a low-dimensional space, with each channel corresponding to an articulator in the vocal tract. This physical embodiment of speech allows for natural and intuitive control and interpretation.
In future work, we will scale the Articulatory Encodec framework to incorporate expressive speech and singing. Also, we will improve the robustness of the system under noisy environments. These efforts will maximize the promising utilities of the articulatory coding of speech.
@@ A.5. Evidence for Excluding Coronal Axes of EMA
In our aritculatory encoding, we largely ignore the coronal axes of EMA data. Here, we provide evidence for such exclusion with a probing analysis. The same linear probing is conducted using SSL features (WavLM) and the original 12 midsagittal articulatory channels. We use EMA-MAE [90] for this experiment since the dataset includes two coronal channels: the left corner of the mouth (LC) and the lateral tongue (LT). Both for LC and LT, the linear probing using SSL features is significantly worse than using articulatory features (for LC, SSL features: $0.68\pm 0.19$, articulatory features: $0.73\pm 0.20$; for LT, SSL features: $0.67\pm 0.19$, articulatory features: $0.72\pm 0.17$). This suggests that those coronal features can be predicted with a fair correlation (0.72-0.73) by a linear combination of the existing 12 channels. These results suggest that the coronal axes can be safely ruled out for the proposed encodec framework.

### v3

@@ VI Conclusion
We propose a novel encoding-decoding framework of speech, SPARC, which mimics the biophysical interface of speech production. Through large-scale training, our proposed SPARC framework achieves a high-performance articulatory inversion and synthesis that can generalize to unseen speakers, with a minimal loss of information compared to ground truth speech. To our knowledge, this is the first demonstration of universal articulatory synthesis that can scale up to an indefinite number of speakers. This universality is supported by a novel speaker identity encoding that embeds highly discriminable speaker-specific voice textures. Additionally, the encoded speaker embedding is effectively disentangled from articulatory features and can be used for accent-preserving voice conversion. Using SPARC, speech can be represented in a low-dimensional space, with each channel corresponding to an articulator in the vocal tract. This physical embodiment of speech allows for natural and intuitive control and interpretation.
In future work, we will scale the SPARC framework to incorporate expressive speech and singing. Also, we will improve the robustness of the system under noisy environments. These efforts will maximize the promising utilities of the articulatory coding of speech.
@@ A.5. Evidence for Excluding Coronal Axes of EMA
In our aritculatory encoding, we largely ignore the coronal axes of EMA data. Here, we provide evidence for such exclusion with a probing analysis. The same linear probing is conducted using SSL features (WavLM) and the original 12 midsagittal articulatory channels. We use EMA-MAE [90] for this experiment since the dataset includes two coronal channels: the left corner of the mouth (LC) and the lateral tongue (LT). Both for LC and LT, the linear probing using SSL features is significantly worse than using articulatory features (for LC, SSL features: $0.68\pm 0.19$, articulatory features: $0.73\pm 0.20$; for LT, SSL features: $0.67\pm 0.19$, articulatory features: $0.72\pm 0.17$). This suggests that those coronal features can be predicted with a fair correlation (0.72-0.73) by a linear combination of the existing 12 channels. These results suggest that the coronal axes can be safely ruled out for the proposed coding framework.